### v1

@@ 1 Introduction
As AI systems are increasingly deployed across diverse linguistic communities, moderation systems [^3] We use the terms ”moderation system”, ”moderation classifier”, and ”guardrail” interchangeably to refer to text filters that assess content safety. are evolving to offer broader multilingual support OpenAI (2024); Meta (2025). However, their effectiveness in localised, low-resource, or code-mixed settings remains limited - leaving critical safety gaps. For instance, multilingual adversarial prompts have been shown to bypass robust filters Yong et al. (2024); Shen et al. (2024); Wang et al. (2024). Singapore exemplifies the challenge: everyday discourse blends English, Chinese, Malay, and Tamil in code-mixed forms like Singlish, featuring local slang, abbreviations, and dialectal variants (Figure 1). Moderation systems that ignore such linguistic nuance risk degraded performance and exploitation in real-world deployments.
@@ 2.1 Multilingual Content Moderation
While commercial moderation APIs offer multilingual support, their efficacy on low-resource or code-mixed languages is often unclear. Open-source models such as LlamaGuard Meta (2024b, 2025), DuoGuard Deng et al. (2025), and PolyGuard Kumar et al. (2025) provide broader coverage, but do not address cultural localisation, limiting their robustness in real-world multilingual environments.
@@ 3 System Overview
Figure 6 demonstrates LionGuard 2 acting as an chatbot guardrail. LionGuard 2 flags localised unsafe content that bypasses both the LLM’s internal safety alignment and commercial moderation classifiers (i.e. OpenAI Moderation). In this example, Singapore-specific acronyms and slang are used to elicit unsafe content from the LLM.
@@ 4.1.3 Automated Labelling with Human Supervision
To mimic low-resource constraints, we employ LLMs to annotate the training examples, but employ statistical methods to ensure as much alignment as possible with human labellers. We begin with a panel of six humans and six LLMs 444o3-mini-low OpenAI (2025), Gemini 2.0 Flash Google (2025), Claude 3.5 Haiku Anthropic (2024), Llama 3.3 70B Meta (2024a), Mistral Small 3 Mistral (2025), and AWS Nova Lite Intelligence (2024). Through the Alt-Test methodology (Calderon et al., 2025), we identified Gemini 2.0 Flash, o3-mini-low, and Claude 3.5 Haiku to be best aligned with our six human annotators, and labeled every example in our dataset with these three selected models (system prompt in Appendix C.5).
@@ 4.2.2 Training a Lightweight Classifier
Table 3: Binary $F_{1}$ scores on Singapore-localised benchmarks. The best results for each dataset are made bold. (–) indicates that the model does not support that language.
@@ 5.3 Red-Teaming by Native Speakers
Employing native annotators, we curated a multilingual test set of 391 cases and benchmarked LionGuard 2 against five other safety models (see details in Appendix E.2). LionGuard 2 surpasses its nearest competitor by 2.4% in Chinese $F_{1}$ and 8.3% in Malay $F_{1}$. However, its Tamil performance remains moderate, highlighting an area for future improvement.
@@ 9 Acknowledgments
Appendix A Demonstration of LionGuard 2 applied as a chatbot guardrail
A demonstration of LionGuard 2 working as a localised content moderator is shown in Figure 5 and Figure 6.
@@ Appendix B Taxonomy
Table 6: Safety Taxonomy: A single text can belong to multiple categories, or none. With the exception of insults and physical violence, severity levels are also available, with Level 2 being more severe than Level 1.
@@ Appendix C Data
C.1 Prompt template for synthetic chat queries
@@ C.2 English datasets used in experiments
PH titles77footnotemark: 7 Used 1M adult-site video titles.
@@
HateXplain (Mathew et al., 2022) Dropped 25,000 English comments with {hate, offensive, neutral}. Only target groups relevant to Singapore retained for experiments.
@@ C.3 Experiments with LLM-translated data
To test whether synthetic Malay/Tamil data could close the low-resource gap, we translated the Singlish corpus with gpt-4o-mini and ran the following training variants: Baseline (final LG2) - 85% Singlish, 15% English; SS-only - Singlish data; SS+MS+TA - Singlish, Malay, and Tamil translated data; and TA-only Tamil translated data.
@@
Table 8: Binary $F_{1}$ on Tamil splits when adding machine-translated data.
@@ C.6 Data Filtering Details
Category re-balancing. Majority of the harms were systematically down-sampled to ensure a more equal distribution of the six major harm categories in the training dataset (Figure  10).
@@
Near-duplicate removal. Using OpenAI’s text-embedding-3-large vectors, we run $k$-nearest neighbors ($k$-NN) and deduplicate pairs above ${}^{\text{th}}$ percentile cosine similarity (Figure 11).
@@ E.2 Red-Teaming by Native Speakers
Model Chinese Malay Tamil Overall
Acc. F1 Acc. F1 Acc. F1 Acc. F1
LionGuard 2 85.5 85.0 79.5 81.4 40.7 41.1 70.6 72.7
AWS Bedrock Guardrails 22.2 17.1 30.7 32.1 35.2 25.7 29.4 25.7
Azure AI Content Safety 81.2 82.6 70.5 73.1 66.7 25.7 72.6 74.2
GCP Model Armor 59.8 65.1 65.7 69.5 62.0 62.9 62.9 66.1
OpenAI Moderation API 53.8 58.8 39.8 44.5 36.1 22.4 43.0 44.4
LlamaGuard 4 12B 44.4 48.2 53.6 59.5 57.4 58.6 51.9 56.0
@@
Table 14: Per-category F1 (in %) on RabakBench. "-" marks unsupported or zero-positive categories.
@@
Table 15: Per-category F1 (in %) on BeaverTails_330k_test.
@@
Table 16: Per-category F1 (in %) on SimpleSafetyTests.

### v2

@@ 1 Introduction
As AI systems are increasingly deployed across diverse linguistic communities, moderation systems [^3] We use the terms ”moderation system”, ”moderation classifier”, and ”guardrail” interchangeably to refer to text filters that assess content safety. are evolving to offer broader multilingual support OpenAI (2024); Meta (2025). However, their effectiveness in localised, low-resource, or code-mixed settings remains limited, leaving critical safety gaps. For instance, multilingual adversarial prompts have been shown to bypass robust filters Yong et al. (2024); Shen et al. (2024); Wang et al. (2024). Singapore exemplifies the challenge: everyday discourse blends English, Chinese, Malay, and Tamil in code-mixed forms like Singlish, featuring local slang, abbreviations, and dialectal variants (Figure 1) Foo and Ng (2024). Moderation systems that ignore such linguistic nuance risk degraded performance and exploitation in real-world deployments.
@@ 2.1 Multilingual Content Moderation
Detecting hateful content in multilingual environment is widely studied in recent years Haber et al. (2023); Hee et al. (2024b); Lee et al. (2024); Hee et al. (2024a). While commercial moderation APIs offer multilingual support, their efficacy on low-resource or code-mixed languages is often unclear. Open-source models such as LlamaGuard Meta (2024b, 2025), DuoGuard Deng et al. (2025), and PolyGuard Kumar et al. (2025) provide broader coverage, but do not address cultural localisation, limiting their robustness in real-world multilingual environments Ng and Carley (2025).
@@ 3 System Overview
Figure 7 demonstrates LionGuard 2 acting as an chatbot guardrail. In this example, Singapore-specific acronyms and slang are used to elicit unsafe content from the LLM. LionGuard 2 flags localised unsafe content that bypasses both the LLM’s internal safety alignment and commercial moderation classifiers (i.e. OpenAI Moderation).
@@
Through the AI Guardian platform, we plan to establish an end-to-end MLOps pipeline to continuously monitor performance and adapt the model to evolving local requirements through retraining and benchmarking of new embeddings.
@@ 4.1.3 Automated Labelling with Human Supervision
To mimic low-resource constraints, we employ LLMs to annotate the training examples, but employ statistical methods to ensure as much alignment as possible with human labellers. We begin with a panel of six humans and six LLMs 444o3-mini-low OpenAI (2025), Gemini 2.0 Flash Google (2025), Claude 3.5 Haiku Anthropic (2024), Llama 3.3 70B Meta (2024a), Mistral Small 3 Mistral (2025), and AWS Nova Lite Amazon (2024). Through the Alt-Test methodology (Calderon et al., 2025), we identified Gemini 2.0 Flash, o3-mini-low, and Claude 3.5 Haiku to be best aligned with our six human annotators, and labeled every example in our dataset with these three selected models (system prompt in Appendix C.5).
@@ 4.2.2 Training a Lightweight Classifier
Table 3: Binary $F_{1}$ scores on Singapore-localised benchmarks. The best results for each dataset are bold. (–) indicates that the model does not support that language.
@@ 5.3 Red-Teaming by Native Speakers
Employing native annotators, we curated a multilingual test set of 391 cases and benchmarked LionGuard 2 against five other safety models (see Figure 5 for examples and Appendix E.2 for further details). LionGuard 2 surpasses its nearest competitor by 2.4% in Chinese $F_{1}$ and 8.3% in Malay $F_{1}$. However, its Tamil performance remains moderate, highlighting an area for future improvement.
Figure 5: Examples from the red-teaming process
@@ 9 Acknowledgments
Appendix A LionGuard 2 as a chatbot guardrail
Figure 6 and Figure 7 demonstrates LionGuard 2 working as a localised content moderator.
@@ Appendix B Taxonomy
Table 6: Safety Taxonomy: A text can belong to multiple categories, or none. Severity levels are available for certain categories.
@@ Appendix C Data
C.1 Prompt template for synthetic queries
@@ C.2 English datasets used in experiments
PH titles77footnotemark: 7 Used 1M adult‐site video titles.
@@
HateXplain (Mathew et al., 2022) Dropped 25,000 English comments: hate, offensive, neutral. Only target groups relevant to Singapore used for experiments.
@@ C.3 Experiments with LLM-translated data
To test whether synthetic Malay/Tamil data could close the low-resource gap, we translated the Singlish corpus with gpt-4o-mini and ran the following training variants.
@@
Table 8: Binary $F_{1}$ on Tamil splits when adding machine-translated data. Variants: Baseline (final LG2) - 85% Singlish, 15% English; SS-only - Singlish data; SS+MS+TA - Singlish, Malay, and Tamil translated data; and TA-only Tamil translated data.
@@ C.6 Data Filtering Details
Category re-balancing. Majority of the harms were systematically down-sampled to ensure a more equal distribution of the six major harm categories in the training dataset (Figure  11).
@@
[…]
@@ E.2 Red-Teaming by Native Speakers
Model Chinese Malay Tamil Overall
Acc. F1 Acc. F1 Acc. F1 Acc. F1
LionGuard 2 85.5 85.0 79.5 81.4 40.7 41.1 70.6 72.7
AWS Bedrock Guardrails 22.2 17.1 30.7 32.1 35.2 25.7 29.4 25.7
Azure AI Content Safety 81.2 82.6 70.5 73.1 66.7 25.7 72.6 74.2
GCP Model Armor 59.8 65.1 65.7 69.5 62.0 62.9 62.9 66.1
OpenAI Moderation API 53.8 58.8 39.8 44.5 36.1 22.4 43.0 44.4
LlamaGuard 4 12B 44.4 48.2 53.6 59.5 57.4 58.6 51.9 56.0
@@
Table 14: Per‐category F1 (in %) on RabakBench. "-" marks unsupported or zero‐positive categories.
@@
Table 15: Per‐category F1 (in %) on BeaverTails_330k_test.
@@
Table 16: Per‐category F1 (in %) on SimpleSafetyTests.